Assume the random variable X can take values of -1, 0, 1, 2. The probabilities are 1/2c 3/4c 5/8c 2/16c. As a result, the constant c =

1
Expert's answer
2022-03-29T11:40:36-0400

"P(-1)+P(0)+P(1)+P(2)=1\\
\frac{1}{2c}+\frac{3}{4c}+\frac{5}{8c}+\frac{2}{16c}=1\\
\frac{8}{16c}+\frac{12}{16c}+\frac{10}{16c}+\frac{2}{16c}=1\\
\frac{32}{16c}=1\\
\frac{1}{c}=\frac{16}{32}=\frac{1}{2}\\
\therefore c=2"